OpenStudy (anonymous):

find the reciprocal : 4

OpenStudy (mertsj):

What are reciprocals?

OpenStudy (anonymous):

r u serious ?

OpenStudy (saifoo.khan):

Hey. reciprocals are something like we have to flip the fractions, like a/b reciprocal will be, b/a.

OpenStudy (mertsj):

Too many cooks. Good bye

OpenStudy (anonymous):

adios....i just want some damn help....

OpenStudy (anonymous):

saifoo...whts tutor...can u help me figure this out or at least how to write

OpenStudy (saifoo.khan):

Yes. So in reciprocal we have to flip the fraction. 4/1 is given to you. flip it, 1/4

OpenStudy (mertsj):

The thing you need to know about reciprocals is that when you multiply them you get 1. \[\frac{2}{3}\times\frac{3}{2}=1\]

OpenStudy (mertsj):

You will need this property when you are solving equations.